# Positive and Risky Message Assessment for Music Products

### v1

@@ Abstract
In this work, we propose a novel research problem: assessing positive and risky messages from music products. We first establish a benchmark for multi-angle multi-level music content assessment and then present an effective multi-task prediction model with ordinality-enforcement to solve this problem. Our result shows the proposed method not only significantly outperforms strong task-specific counterparts but can concurrently evaluate multiple aspects.
1 Introduction
@@
The American Academy of Pediatrics (AAP) holds the opinion that parents should be informed of pediatricians’ concerns regarding potentially harmful effects of music lyrics (American Academy of Pediatrics, 1996). Policymakers have also taken action; for instance, the Recording Industry Association of America (RIAA) introduced the Parental Advisory Label (PAL) program in 1985 to identify audio products with potentially inappropriate content. Such labels were created to draw parents’ attention to products that may not be suitable for their children. Recent studies from the NLP community have made advances in automating the content rating process. Chin et al. and Fell et al. presented machine learning-based methods to automatically classify explicit/non-explicit lyrics in different languages. The learning objectives are derived from content rating systems like PAL, and these works treat the problem as a binary classification task. Despite these existing achievements, however, neither the current PAL system nor the corresponding automated approaches are capable of appraising a music product with diverse information about content suitability.
@@
In this work, we introduce a novel NLP task: assessing the positive and risky messages of a music item. We study the messages that a music item conveys from five significant dimensions regarding appropriateness: Positive Messages, Violence, Substance Consumption, Sex, and Consumerism along three degrees of severity. In the light of those perspectives, we propose a multi-task method that successfully incorporates both content aspect correlations and severity ordinalities to better assess music products. Our research focuses specifically on music performed with vocal techniques (singing, rapping, etc). In such a format of music, the lyrics play a dominant part in conveying the message and opinion from the artists, which has a major impact on the listener’s thoughts and minds. With such an automated approach of music content assessment, any music provider will benefit from service quality improvements for young users and their parents by offering better recommendation or content advisory. Meanwhile, lyricists may also leverage the system to evaluate the portrayal of risky behaviors like ‘get drunk and be somebody’ even at their early stage of creation.
Our contribution: To our best knowledge, this work, for the first time, presents an exploration on assessing music items from both positive and risky aspects on multiple levels. We develop an effective benchmark and made the dataset creation script publicly available for the community. We further propose an effective multi-task ordinality-enforced rating method that can jointly assess different risky aspects with state-of-the-art performance.
2 Music rating and lyrics
@@
To build up a reliable benchmark for assessing positive and risky messages in such lyrics, we exploit expert ratings provided by Common Sense Media (CSM). CSM is a non-profit organization caring for kids’ digital well-being. It hosts a media rating platform that is supported by childhood development experts. The rating system covers various aspects of childhood development and age appropriateness. In this work, based on the focal points from Youth Risk Behavior Surveillance System (YRBSS) Kolbe et al. (1993) introduced by the USA Centers for Disease Control and Prevention, we choose three explicitly risky aspects Violence, Substance Consumption, and Sex. We also pick Consumerism that relates to drawing interest in the acquisition of goods, often associated with a “feel good" experience. Because Consumerism potentially affects the emotional health and identity development in youth Hill (2011). Additionally, we also study Positive Messages in music products in view of the fact that positive values and behaviors are always encouraged in adolescent development.
2.1 Dataset facts and analysis
@@
With CSM expert ratings, we developed a dataset based on publicly available lyrics from the Internet. We collected 1,119 music items (consisting of 10,661 songs) including standard albums, extended plays (EP), long plays (LP), and CD singles. We refer to the collections of songs as albums in this work for simplification. The partition of CD singles and the albums are shown in Table 1. The age recommendation spans from 2 to 18+.
@@
Figure 1: Overall caption for both images
@@
We gained permission from CSM to use the music rating data for research, however, by the time we submit this work, the music reviews are no longer shown to the public on CSM’s website until further updates. We still release the names of expert-rated products and links to their lyrics studied in this work for the community.
The music rating data comes from Common Sense Media (CSM). Non-member users can only browse up to three expert product reviews for free each month. We gained permission from CSM to use the music rating data for research, however, by the time we submit this work, the music reviews are no longer shown on CSM’s website until further updates. We still release the names of expert-rated products and links to their lyrics studied in this work for the community. For the lyrics, we do not release them directly for copyright reasons, It is very easy to find lyrics through our provided links and search engines. We make our best effort to collect lyrics from the open Internet but we recognize that there are still very few songs missing.
2.1.1 Inter-dimensional correlation analysis
@@
We can easily observe from the heat map that typical physical risky behaviors Violence, Substance Consumption, and Sex have a positive correlation with each other, while Consumerism has a positive correlation with those three but less strong. It is no surprise that Positive Messages has a significant negative correlation with three physical risky behaviors. This interdimensional behavior intuitively inspire us to leverage such correlations to design relevant machine learning strategies.
2.1.2 Positive/Risky behavior v.s. Explicitness
In the context of music, explicitness often refers to the content that contains strong language or violence, sex, or substance abuse depictions (from RIAA). It is a generalized description that has a high coincidence with the risky message we studied in this work. Since it is not practical to collect gold labels of explicitness for every music item we studied, we apply an explicit lyrics classifier trained on 438k English lyrics from a previous work Fell et al. (2020). We can further explore the correlation between the explicitness probability and the level of positive and risky messages. Figure 3 illustrates the pattern of explicitness probability towards severity levels of different aspects.
@@
It is evident that the message ratings are nicely correlated with the explicitness values. For all of the risky message aspects, the explicitness raises as the message ratings go from low to high. Interestingly, the most abstract message, Consumerism, also follows this pattern. On the contrary, Positive Messages also gives a clear negative correlation against explicitness. This correlation study informs us the message ratings for different aspects can be a good complement or alternative to show the inappropriateness for the music items but are more informative, which affirms the significance of our study.
@@ 3 Methodology
Emotion-guided twin model: We propose to encode the text from two perspectives: universal semantic representation and emotion-centered semantic representation. We employ two pretrained Transformer models finetuned with semantic textual similarity (Sentence-BERT Reimers and Gurevych (2019)) and emotion detection (fine-tuned Distilled RoBERTa on emotion detection tasks Hartmann (2022)) respectively. The final representation is the concatenation of these two types of embeddings.
Joint prediction over multiple aspects: From the dataset study, we found different aspects, specifically risky behaviors, have a potential correlation with each other. For instance, the presence of violence in a music item might appear concurrently with lyrics depicting substance use. The learning objective is a joint loss of predicting ratings of multiple aspects. To fortify the representation uniqueness for different aspects, we further design an aspect-aware attention module to learn specialized features for each aspect. The module begins with a projection layer with non-linearity and further learns weights through an aspect-differentiation matrix. The final representation $\boldsymbol{x}_{out}$ comes from the addition of input $\boldsymbol{x}_{in}$ with residual over skip connections as described in Equation 1.
Ordinality-enforcement techniques: The severity ratings are discrete interclass ordinal variables instead of independent categorical labels. Typical classification models usually ignore such correlations between ordinal categories. In this work we apply three ordinality-enforcement techniques that are better suited for the task:
@@
Binary attributes transformation Frank and Hall (2001): This method first divide the sorted ordinal label set (with $n$ elements) into two sets by splitting at every two adjacent elements. It is obvious that there are $n-1$ ways of division. Then we further treat each division case as a binary classification problem and predict the ordinal value on the corresponding set. This method applies more than one classifier to exploit the ordering information several times. In the setting of this problem, we split the 3-class ordinal levels for each aspect into two cases considering possible probability intervals:
$\operatorname{Pr}(y_{i}\leqslant\text{Low}),\operatorname{Pr}(y_{i}>\text{Low})\;$ (3)
Therefore, we apply two binary classifiers to those two scenarios.
Soft label Diaz and Marathe (2019): This method introduces a label softening function to convert ordinal category values as discrete distribution. A class label $y_{i}$ from label set is encoded using the following formula into a soft label $c_{i}^{soft}$.
The Kullback-Leibler divergence is used as loss function to measure the difference between the predicted probability and the soft label. $\phi\left(y_{t},y_{i}\right)=\left|y_{t}-y_{i}\right|$ is chosen as the metric penalty for the sake of simplicity.
All of the selected ordinality-enforcement techniques are applied respectively to the base model with a multi-tasking module.
4 Experiments and Results
@@
Table 4: An input perturbation study on the behavior of the proposed ranking-classification model. We choose three risky aspects - Violence, Substance consumption, and Sex - as it performs the best. The numbers indicate the absolute probability change of the original prediction result. A double down arrow $\Downarrow$ indicates the predicted severity downgraded by two levels, a single down arrow $\downarrow$ means downgraded by one, and an up arrow $\uparrow$ represents upgraded by one.
[…]
5 Discussion and analysis
Speaking of aspects, three types of explicit content, Violence, Substance, and Sex are relatively easier for the model to predict than Consumerism and Positive messages. The reason could be that the latter aspects are more abstract concepts comparing to risky behaviors that can be described in the lyrics with an overt manner. For Consumerism, we hypothesize a model that can better capture signals on ’goods’ and encouragement for acquisition may work more effectively. Unlike explicit questionable content like Violence, Positive messages is even harder to intuitively find a clear textual pattern since positive is more heterogeneous. We may need the model to gain a high-level understanding of the content to distinguish the quantity of positive value that a song or an album wants to deliver.
5.1 Ablation study
Ablation Perf change
Best performing model 0.00
Aspect-aware module -0.30
Emotion-guided twin model -0.11
Ordinality enforcement (binary) -0.24
Multitask joint prediction 1.17
Backbone only + multitask -2.52
5.2 Saliency analysis
@@
5.3 Error analysis
Figure 5 shows the confusion matrix of the prediction results of the best-performing model on one-fold. In general, the proposed method can capture the ordinal information well because wrong predictions that cross two levels (predict low to high or high to low) are rare. Specifically, the model struggled in giving correct predictions on high Consumerism. We suppose the number of training instances of high Consumerism is relatively small. The same case happened to high Positive predictions. We hypothesize the heterogeneous nature of Positive content makes it challenging to predict.
Figure 5: Prediction confusion matrix of the best-performing method on all 5 aspects.
5.4 Case study of unsuccessful predictions
@@
    We suppose the density of such obvious expressions is low, but many of the songs are expressing a sense of spirit in pursuit of happiness and love. That makes this album has a medium rating in Positive but hard for the model to capture.
  • Album: A Hard Day’s Night (1964) by The Beatles: The model gives a critical wrong prediction on Positive: predict as low while the ground truth is high. A majority of songs in this album express a deep and genuine emotion about love, such as:
@@
    The reason that the model made such a mistake could be the lyrics in this album looks plain but the overall expression is obviously positive and constructive.
  • CD single: Labels or Love (2008) by Fergie: The model gives Sex medium, Substance consumption medium, and Consumerism low, while the correct answer are Sex low, Substance consumption low, and Consumerism high. For Sex medium and Substance consumption, the false positive may come from the lexical signals such as sexy, kiss and bag. For Consumerism, there are not only expressions to the encouragement of buying goods:
@@
    We suppose the model didn’t take a direct narrative of shopping as a strong signal of Consumerism. Meanwhile, the names of the luxury brand are rare words in the corpus and lack supervision signals. So the model fell hard on this prediction.
6 Ethical considerations and limitations

### v2

@@ Abstract
In this work, we introduce a pioneering research challenge: evaluating positive and potentially harmful messages within music products. We initiate by setting a multi-faceted, multi-task benchmark for music content assessment. Subsequently, we introduce an efficient multi-task predictive model fortified with ordinality-enforcement to address this challenge. Our findings reveal that the proposed method not only significantly outperforms robust task-specific alternatives but also possesses the capability to assess multiple aspects simultaneously. Furthermore, through detailed case studies, where we employed Large Language Models (LLMs) as surrogates for content assessment, we provide valuable insights to inform and guide future research on this topic. The code for dataset creation and model implementation is publicly available at https://github.com/RiTUAL-UH/music-message-assessment.
Keywords: Document Classification, Text Categorization, Text Mining, Tools, Systems, Applications
[…]
Positive and Risky Message Assessment for Music Products
Yigeng Zhang1, Mahsa Shafaei1, Fabio A. González2, Thamar Solorio1,3
1University of Houston, Houston, USA
2Universidad Nacional de Colombia, Bogotá, Colombia
3MBZUAI, Masdar City, United Arab Emirates
1{yzhang168,mshafaei,tsolorio}@uh.edu
2fagonzalezo@unal.edu.co
Abstract content
1.   Introduction
@@
The American Academy of Pediatrics (AAP) holds the opinion that parents should be informed of pediatricians’ concerns regarding the potentially harmful effects of music lyrics (American Academy of Pediatrics, 1996). Policymakers have also taken action; for instance, the Recording Industry Association of America (RIAA) introduced the Parental Advisory Label (PAL) program in 1985 to identify audio products with potentially inappropriate content. Such labels were created to draw parents’ attention to products that may not be suitable for their children. Recent studies from the NLP community have made advances in automating the content rating process. Chin et al. and Fell et al. presented machine learning-based methods to automatically classify explicit/non-explicit lyrics in different languages. Further work Rospocher and Eksir (2023) studied detecting explicitness in several aspects such as strong language and language that refers to violence. The learning objectives are derived from content rating systems like PAL, and these works treat the problem as a binary classification task. Despite these existing achievements, however, neither the current PAL system nor the corresponding automated approaches are capable of appraising a music product with severity information about content suitability.
@@
In this work, we introduce a novel NLP task: assessing the positive and risky messages of a music item. We study the messages that a music item conveys from five significant dimensions regarding appropriateness: Positive Messages, Violence, Substance Consumption, Sex, and Consumerism along three degrees of severity. In light of those perspectives, we propose a multi-task method that successfully incorporates both content aspect correlations and severity ordinalities to better assess music products. Our research focuses specifically on music performed with vocal techniques (singing, rapping, etc). In such a format of music, the lyrics play a dominant part in conveying the message and opinion from the artists, which has a major impact on the listener’s thoughts and minds. Leveraging an automated approach for music content assessment can offer numerous advantages to various stakeholders in the music industry. For music providers, it promises enhanced service quality, particularly benefiting younger listeners and their guardians by facilitating better content recommendations and advisories. Furthermore, lyricists can utilize this system during the early stages of their creative process to assess the portrayal of potentially risky behaviors, such as ‘get drunk and be somebody’. Additionally, the automated identification of questionable content can expedite the creation of clean lyric versions, enabling artists to produce safe renditions more efficiently.
Our contribution: To the best of our knowledge, this study represents the first exploration into assessing music items, evaluating both their positive and risky dimensions across multiple levels. We have established a comprehensive benchmark and have made the dataset creation method accessible to the wider research community [^1] https://github.com/RiTUAL-UH/music-message-assessment.. Additionally, we introduce an effective multi-task, ordinality-enforced rating approach that jointly assesses diverse risk factors, delivering state-of-the-art results. As part of our thorough analysis, we also evaluate the efficacy of Large Language Models (LLMs) and provide in-depth discussions, paving the way for future research for this topic.
2.   Music rating and lyrics
@@
To build up a reliable benchmark for assessing positive and risky messages in such lyrics, we exploit expert ratings provided by Common Sense Media (CSM) [^2] https://www.commonsensemedia.org.. CSM is a non-profit organization caring for kids’ digital well-being. It hosts a media rating platform that is supported by childhood development experts. The rating system covers various aspects of childhood development and age appropriateness. In this work, based on the focal points from Youth Risk Behavior Surveillance System (YRBSS) Kolbe et al. (1993) introduced by the USA Centers for Disease Control and Prevention, we choose three explicitly risky aspects Violence, Substance Consumption, and Sex. We also pick Consumerism that relates to drawing interest in the acquisition of goods, often associated with a “feel good" experience. Because Consumerism potentially affects the emotional health and identity development in youth Hill (2011). Our dataset also includes a category dedicated to Positive Messages within music products. The automatic detection of such uplifting content offers numerous valuable applications, enhancing user experience. This could serve as a counter to the often conflicting messages found in mainstream lyrics. Furthermore, it might be leveraged as a form of supportive therapy to positively influence users’ moods.
2.1.   Dataset facts and analysis
@@
With CSM expert ratings, we developed a dataset based on publicly available lyrics from the Internet. We collected 1,119 music items (consisting of 10,661 songs) including standard albums, extended plays (EP), long plays (LP), and CD singles. We refer to the collections of songs as albums in this work for simplification. The partition of CD singles and the albums are shown in Table 1. The age recommendation spans from 2 to 18+. Notably, the music items we consider in this work are in English and comprise a heavily Western-centric dataset.
@@
Figure 1: Lyrics length distribution for CD singles and albums.
@@
The music rating data comes from Common Sense Media (CSM). Non-member users can only browse up to three expert product reviews for free each month. We gained permission from CSM to use the music rating data for research, however, by the time we submit this work, the music reviews are no longer shown on CSM’s website until further updates. We still release the names of expert-rated products and links to their lyrics studied in this work for the community. We do not release the lyrics directly due to copyright considerations; however, lyrics can be easily accessed through the links we provide and via search engines. While we’ve made every effort to collect lyrics from the open Internet, we acknowledge that some songs might be missing from certain albums.
2.1.1.   Inter-dimensional correlation analysis
@@
We can easily observe from the heat map that typical physical risky behaviors Violence, Substance Consumption, and Sex have a positive correlation with each other, while Consumerism has a positive correlation with those three but less strong. It is no surprise that Positive Messages has a significant negative correlation with three physical risky behaviors. This interdimensional behavior intuitively inspires us to leverage such correlations to design relevant machine learning strategies.
2.1.2.   Positive/Risky behavior v.s. Explicitness
In the context of music, explicitness often refers to content that contains strong language or violence, sex, or substance abuse depictions (from RIAA). It is a generalized description that has a high coincidence with the risky message we studied in this work. Since it is not practical to collect gold labels of explicitness for every music item we studied, we apply an explicit lyrics classifier trained on 438k English lyrics from a previous work Fell et al. (2020). We can further explore the correlation between the explicitness probability and the level of positive and risky messages. Figure 3 illustrates the pattern of explicitness probability towards severity levels of different aspects.
@@
The message ratings consistently correlate with explicitness values. For all risky message aspects, explicitness increases as ratings transition from low to high. Even the abstract message of Consumerism aligns with this trend. In contrast, Positive Messages exhibit a distinct negative correlation with explicitness. These correlations suggest that message ratings can serve as a valuable complement or alternative to traditional metrics of music appropriateness. They provide a richer perspective and highlight the significance of our study.
@@ 3.   Methodology
Emotion-guided twin model: We propose to encode the text from two perspectives: general-purpose semantic representation and emotion-centered semantic representation, as emotion information has been proved to be an effective complement for various language understanding tasks Shafaei et al. (2020); Samghabadi et al. (2020). We employ two pretrained Transformer models finetuned with semantic textual similarity (Sentence-BERT Reimers and Gurevych (2019)) and emotion detection (task-specific Distilled RoBERTa on emotion detection Hartmann (2022)) respectively. The final representation is the concatenation of these two types of embeddings.
Joint prediction over multiple aspects: From the dataset study, we found different types of positive and risky messages, specifically questionable behaviors, have a potential correlation with each other. For instance, the presence of violence in a music item might appear concurrently with lyrics depicting substance use. The learning objective is a joint loss of predicting ratings of multiple aspects. To fortify the representation uniqueness for different aspects, we further design an aspect-aware attention module to learn specialized features for each aspect. The module begins with a projection layer with non-linearity and further learns weights through an aspect-differentiation matrix. The final representation $\boldsymbol{x}_{out}$ comes from the addition of input $\boldsymbol{x}_{in}$ with residual over skip connections as described in Equation 1.
Ordinality-enforcement techniques: The severity ratings are discrete interclass ordinal variables instead of independent categorical labels. Typical classification models usually ignore such correlations between ordinal categories. In this work, we apply three ordinality-enforcement techniques that are better suited for the task. All of the ordinality-enforcement techniques are applied respectively to the base model with a multi-tasking module.
[…]
  • Binary attributes transformation Frank and Hall (2001): This method tackles ordinal regression by dividing the sorted ordinal label set, containing $n$ elements, into two subsets at every possible pair of adjacent elements. This results in $n-1$ potential splits. Each split transforms the ordinal regression problem into a binary classification task, where the goal is to predict whether the ordinal value $y_{i}$ falls before or after the split point within the set. This approach applies multiple classifiers to leverage the ordinal information repeatedly. Specifically, in the setting of this problem with three ordinal classes (low, medium, and high), we consider two binary splits: one between low and medium, and another between medium and high. This results in two binary classifications:
    $\operatorname{Pr}(y_{i}\leqslant\text{Low}),\operatorname{Pr}(y_{i}>\text{Low});$ (3)
    In this setting, for each classification, we apply a binary classifier to predict which split the prediction will fall in, utilizing the ordinal information multiple times aiming to improve the prediction performance.
  • Soft label Diaz and Marathe (2019): This method introduces a label softening function to convert ordinal category values into a probability distribution across categories. A class label $y_{i}$ from the label set is encoded into a soft label $c_{i}^{soft}$ using the following formula for a specific true rating rank $y_{t}$.
    The Kullback-Leibler divergence is used as the loss function to measure the difference between the predicted probability and the soft label. $\phi\left(y_{t},y_{i}\right)=\left|y_{t}-y_{i}\right|$ is chosen as the metric penalty for the sake of simplicity.
@@
4.   Experiments and Results
To evaluate the effectiveness of our proposed method in music content assessment, we benchmarked several popularly used classification methods and models from related works in media rating. We choose TF-IDF and Bag-of-Word-Vectors (Averaged GloVe Pennington et al. (2014) embeddings) with SVM classifiers, TextCNN Kim (2014) and TextRCNN Lai et al. (2015). We experiment with three deep models that are designed for media rating problems:
  1. 1.
    An RNN model with attention for predicting movie MPAA ratings based on movie dialogue scripts Shafaei et al. (2020) as lyrics are also sequences of utterances;
  2. 2.
    A BERT model Devlin et al. (2019) as used for classifying explicitness in Fell et al. (2020);
  3. 3.
    An RNN-Transformer backbone model of state-of-the-art in rating severity for age-restricted content in movies Zhang et al. (2021).
The training-development-test split uses an 80/10/10 ratio with data shuffled. We choose macro F1 as the classification performance metric because the label distribution from the dataset is highly imbalanced. Experimental results are shown in Table 3.
@@ 4.   Experiments and Results
5.   Discussion and analysis
Speaking of aspects, three types of explicit content, Violence, Substance, and Sex are relatively easier for the model to predict than Consumerism and Positive messages. The reason could be that the latter aspects are more abstract concepts compared to risky behaviors that can be described in the lyrics in an overt manner. For Consumerism, we hypothesize that a model which can more effectively capture signals related to ’goods’ and the promotion of purchases might perform better. Unlike explicit questionable content like Violence, Positive messages is even harder to intuitively find a clear textual pattern since positive is more heterogeneous than the other aspects. We may need the model to gain a high-level understanding of the content to distinguish the quantity of positive value that a song or an album delivers.
5.1.   Ablation study
Figure 5: Prediction confusion matrix of the best-performing method on all 5 aspects. The x-axis indicates the predicted values and the y-axis indicates the ground-truth labels.
Ablation Perf change
Best performing model 57.88
Aspect-aware module 57.58 (-0.30)
Emotion-guided twin model 57.77 (-0.11)
Ordinality enforcement (binary) 57.53 (-0.24)
Multitask joint prediction 59.05 (1.17)
Backbone only + multitask 55.36 (-2.52)
@@ 5.1.   Ablation study
Table 5: An input perturbation study on the behavior of the proposed ranking-classification model. We choose three risky aspects - Violence, Substance consumption, and Sex - as this model yields the best performance. The numbers indicate the absolute probability change of the original prediction result. A double down arrow $\Downarrow$ indicates the predicted severity downgraded by two levels, a single down arrow $\downarrow$ means downgraded by one, and an up arrow $\uparrow$ represents upgraded by one.
5.2.   Saliency analysis
@@
5.3.   Error analysis
Figure 5 shows the confusion matrix of the prediction results of the best-performing model from a single fold of the cross-validation. In general, the proposed method can capture the ordinal information well because wrong predictions that cross two levels (predict low to high or high to low) are rare. Specifically, the model struggled to give correct predictions on high Consumerism. We suppose the number of training instances of high Consumerism is relatively small. The same case happened to high Positive predictions. We hypothesize the heterogeneous nature of Positive content makes it challenging to predict.
Figure 6: Prediction confusion matrix of the best-performing LLM method on all 5 aspects. The x-axis indicates the predicted values and the y-axis indicates the ground-truth labels.
5.4.   Case study: unsuccessful predictions
@@
    We suspect that explicit expressions of positivity as in the example are sparse in the songs. However, many songs in the album convey a spirit of pursuing happiness and love. This nuance gives the album a medium rating in Positive, but it is challenging for the model to capture.
  • Album: A Hard Day’s Night (1964) by The Beatles: The model gives a critical wrong prediction on Positive: predict as low while the ground truth is high. The majority of the songs in this album express deep and genuine emotions about love, with lyrics like:
@@
    One possible reason the model failed is that the lyrics do not contain words that explicitly convey strong positive signals of companionship. However, the overall sentiment is clearly positive and constructive.
  • CD single: Labels or Love (2008) by Fergie: The model gives Sex medium, Substance consumption medium, and Consumerism low, while the correct answer are Sex low, Substance consumption low, and Consumerism high. For Sex medium and Substance consumption, the false positive may come from lexical signals such as sexy, kiss and bag. For Consumerism, there are not only expressions to the encouragement of buying goods:
@@
    We suspect that the model did not interpret a direct narrative about shopping as a strong indicator of Consumerism. Additionally, the names of luxury brands might be rare in the corpus, leading to a lack of supervision signals. As a result, the model struggled with this prediction.
Vio Sub Sex Con Pos Avg
Simple context 54.91 49.87 49.09 45.35 34.35 46.71
Rich context 53.51 54.06 53.79 40.94 35.33 47.53
Simple context + CoT 53.72 37.83 43.22 45.95 29.77 42.10
Rich context + CoT 55.52 50.18 54.03 40.82 29.05 45.92
Table 6: Zero-shot evaluation results from gpt-3.5-turbo. Only CD-singles are evaluated due to the intrinsic token size limitations of this LLM.
5.5.   Case study: LLM as content judges
Recent advancements in Large Language Models (LLMs) have showcased impressive natural language understanding and adaptability across a multitude of tasks. Motivated by these advancements, our study aims to explore the potential of LLMs in assessing content within music products. Specifically, we leverage the gpt-3.5-turbo API OpenAI (2022); Ouyang et al. (2022) as surrogate evaluators. Our primary focus is on rating five specific aspects of the content. The central hypothesis of this study is that, despite their inherent limitations and lack of access to a supervision signal, LLMs can provide content assessments that are both meaningful and comparable in accuracy to other deep learning methods. Due to the context length constraints of LLMs, we limit our evaluation to CD-singles, setting a token cap of 3000 to ensure the model’s efficient functioning.
The experiments were structured in three distinct formats:
  • Simple Context: The LLM is directed to rate each song across the five aspects without any supplementary information.
  • Rich Context: Before prompting the LLM, a detailed description of the five aspects is provided in the context.
  • Chain-of-Thought (CoT): Building upon the CoT approach Wei et al. (2022), known for enhancing LLM performance in complex reasoning tasks, we feed the model with exemplar prompts as context and subsequently instruct it to complete the rating task.
Table 6 presents the evaluation results. While LLM-based approaches have their merits, they did not demonstrate remarkable efficiency in this content rating context. Although providing a richer context yielded marginally superior outcomes, no method consistently outperformed the others. Interestingly, LLM assessments aligned with patterns observed in our baseline and proposed methods. Specifically, the LLM found it more straightforward to evaluate the explicit aspects of Violence, Substance, and Sex, but faced challenges with Consumerism and Positive. It’s important to note that this experimental design is simpler than previous sections, given it exclusively assesses CD-singles rather than an assorted selection from an album.
Further analysis of the LLM’s performance was conducted by examining the confusion matrix of the top-performing model, as depicted in Figure 6. The model displayed some notable patterns in its behavior. Specifically:
  • For categories like Violence, Sex, and Consumerism, the model tended to underestimate their respective severities.
  • When rating Substance, the model frequently struggled to give medium presence ratings.
  • In contrast, the evaluation of Positive content often resulted in an overestimation of a song’s positive messages, different from patterns observed in prior experiments of baseline and proposed methods.
It’s essential to recognize the intrinsic limitations of the LLM. We could not apply the same training and assessment methods to the LLM as we did in previous sections. Although CD singles are a subset of the broader music product collection, they retain a consistent data property. While our comparison does not strictly align with traditional comparative analysis standards, due to potential disparities in data distribution and features across datasets, it nonetheless provides valuable insights. These insights can guide model benchmarking and optimization, even if not strictly empirical. Our decision to evaluate closed models like gpt-3.5-turbo—which is non-reproducible due to its proprietary nature—stems from a desire to explore the capabilities of such models. We urge readers to interpret these particular results as exploratory, rather than as fixed benchmarks.
While LLMs have demonstrated proficiency in a variety of NLP tasks, their performance in our specialized context of content assessment was not on par. This discrepancy is understandable given that these versatile models are not trained for such tasks. Consequently, their judgments might not always align with the expert opinions of professionals in media research and childhood development. Recognizing this, our forthcoming research aims to explore the development and analysis of task-specific LLMs for content safety. We are optimistic that such an approach will yield meaningful insights and enhanced performance.
6.   Conclusion
In this paper, we introduce a novel task to the NLP community: predicting the intensity of various aspects of music, spanning from objectionable content to positive messages. By analyzing music product lyrics, we investigate multiple dimensions of messages conveyed to listeners. Our research problem and approach are intended to foster deeper investigations into music content assessment. The multi-task ordinality-enforcement model we present has shown promising effectiveness for this type of challenge with ordinal properties. The case studies, along with our exploration using Large Language Models (LLMs) as surrogate evaluators, highlight the inherent complexities of the message assessment problem, calling for the need for continued community engagement and research.
Ethical considerations and limitations